OpenStudy (anonymous):

Is the ordered pair a solution. Y=4x+9;(-2,1)

OpenStudy (anonymous):

Check. Let y=1, and x=-2. 1=4(-2)+9? 1=-8+9? 1=1 yes

OpenStudy (texaschic101):

nicely done larryboxaplenty :)